Rain Drop 2 A spherical rain drop carries a charge of 8.70E-8 C and has a potential of 1.82E+5 V at the surface. Find the surface area of the drop assuming that the potential equals zero at infinity Submit Answer Tries 0/10
Answer #1

Potential at the surface of a sphere is given by:

V = k*Q/R

V = Potential = 1.82*10^5 V

k = electrostatic constant = 9*10^9

Q = Charge on sphere = 8.70*10^-8 C

R = radius of sphere = ?

Using known values:

R = k*Q/V

R = 9*10^9*8.70*10^-8/(1.82*10^5)

R = 4.30*10^-3 m

Now Surface area of spherical drop is given by:

S = 4*pi*R^2

S = 4*pi*(4.30*10^-3)^2

S = 2.32*10^-4 m^2

S = 2.32E-4 m^2

Please Upvote.

Comment below if you have any query.